OpenStudy (anonymous):

A cylinder and a rectangular prism have the same volume and same height. The base of the prism is a square with side length 9 cm. What is the approximate radius of the cylinder? A. 4.5 cm B. 5.1 cm C. 12.9 cm D. 25.8 cm

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (campbell_st):

well the area of the square base of the prism is \[A = 9^2\] and given the circular base of the cylinder has the same area you have \[9^2 = \pi r^2\] solve for r. hope it helps

OpenStudy (anonymous):

so the answer is @campbell_st

OpenStudy (campbell_st):

i don't know... I just explained how you can find it

OpenStudy (anonymous):

@johnweldon1993 plz help

OpenStudy (anonymous):

is it b 5.1 @campbell_st

OpenStudy (campbell_st):

well \[r = \sqrt{\frac{81}{\pi}}\]